# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v1

@@ 4.1 Main Results
Table 1: Performance comparison using Llama2-7b [2] and GraphTransformer [37] as backbones and same retrieval settings for all methods. The table reports the mean and standard deviation (std) results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report both F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from [18] and were not reproduced due to their poor performance. The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement’ represents the gain of Align-GRAG over the second-best baseline. ‘*’ indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better.
@@
Inference-$\ddagger$
Zero-shot 0.5650 0.3974 - 0.4106 -
Zero-CoT [40] 0.5704 0.5260 - 0.5130 -
CoT-BAG [41] 0.5794 0.5680 - 0.3960 -
KAPING [42] 0.6227 0.4375 - 0.5264 -
@@
gte-base [22] 0.8557$\pm$0.0144 0.8556$\pm$0.0095 0.5378$\pm$0.0044 0.7373$\pm$0.0064 0.5251$\pm$0.0052
gte-large [22] 0.8776$\pm$0.0095 0.8592$\pm$0.0074 0.5392$\pm$0.0013 0.7340$\pm$0.0044 0.5374$\pm$0.0038
bge-reranker-base [21] 0.8534$\pm$0.0159 0.8577$\pm$0.0029 0.5323$\pm$0.0052 0.7397$\pm$0.0012 0.5254$\pm$0.0010
bge-reranker-large [21] 0.8612$\pm$0.0184 0.8644$\pm$0.0060 0.5366$\pm$0.0045 0.7391$\pm$0.0093 0.5401$\pm$0.0077
G-RAG [44] 0.8484$\pm$0.0174 0.8474$\pm$0.0147 0.5181$\pm$0.0023 0.7114$\pm$0.0113 0.5080$\pm$0.0041
G-RAG-RL [44] 0.8478$\pm$0.0112 0.8509$\pm$0.0142 0.5291$\pm$0.0066 0.7167$\pm$0.0039 0.5185$\pm$0.0026
@@
$\text{GRAG}_{PCST}^{\dagger}$ [45] 0.8805$\pm$0.0050 0.8561$\pm$0.0052 0.5355$\pm$0.0049 0.7485$\pm$0.0104 0.5503$\pm$0.0035
Align-GRAG (Ours) 0.8992$\pm$0.0124 0.8804$\pm$0.0106 0.5445$\pm$0.0041 0.7626$\pm$0.0063 0.5700$\pm$0.0039
In this section, we conduct extensive experiments on the ExplaGraphs, SceneGraphs, and WebQSP datasets, comparing the performance of our method with 16 baseline methods. All baselines are implemented using the same retrieved text, and the same backbones, Llama2-7b [2] and GraphTransformer [37]. As shown in Table 1, the overall results demonstrate the effectiveness of Align-GRAG, which achieves performance improvements across all metrics on three datasets. Notably, on WebQSP dataset, Align-GRAG achieves a remarkable 4.76% improvement in Accuracy compared to the second-best method. This improvement can be attributed to the innovative design of Align-GRAG, particularly its alignment-based pruning and optimization strategies for graph structures, showcasing its potential in complex graph reasoning and knowledge-based question answering tasks.
Comparing with Inference-only methods. They rely solely on the reasoning capabilities of LLM without task-specific optimization and, as a result, perform poorly. Specifically, the Zero-shot approach, which does not depend on any carefully constructed prompts, suffers from limited performance due to its inability to leverage task-specific knowledge.
@@
Comparing with reranker-based methods. gte-large achieves the second-best performance on both the ExplaGraphs and WebQSP datasets (F1 metric), outperforming other reranker approaches. However, it still falls significantly short of Align-GRAG. This demonstrates that simply performing node and triple reranking is insufficient to fully capture the graph structural information. Although G-RAG-RL also utilizes a GNN encoder to model graph structures, it rely on similarity with the query as a label, resulting in poor performance. This validates the effectiveness of our approach, which uses reasoning chains summarized by LLM as optimization target. By aligning graph reasoning with LLM-derived reasoning chains, Align-GRAG achieves superior performance, showing that this alignment is crucial for graph-based reasoning tasks.
@@
4.2 Ablation Study
Table 2: Ablation study on different alignment strategy.
Method WebQSP
@@ 4.2 Ablation Study
Align-GRAG 0.5445 0.7626 0.5700
w/o Representation Alignment 0.5458 0.7586 0.5675
w/o Node Alignment 0.5344 0.7371 0.5339
w/o Both 0.5348 0.7328 0.5216
Random Alignment 0.4617 0.6861 0.4865
We perform an ablation study on WebQSP using three metrics to evaluate how the proposed modules contribute to performance improvement. As illustrated in Table 2, we examine the following variants: (1) w/o Node Alignment: Removes the KL loss and knowledge pruning components. (2) w/o Representation Alignment: Excludes the contrastive loss and the graph embedding input to the LLM. (3) w/o Both: Removes both Node Alignment and Representation Alignment. (4) Random Alignment: Optimizes the aligner module using randomly generated labels. We observe that removing the Node Alignment module significantly lowers all evaluation metrics compared to the full model (Align-GRAG), indicating the effectiveness of node-level alignment optimization and node pruning. In contrast, while excluding the Representation Alignment module also leads to a performance decline, the impact is comparatively less significant. Performance further deteriorates when both modules are removed (w/o Both), highlighting the critical role of dual alignment between node and representation in filtering irrelevant knowledge and bridging the representation gap between the graph and language. The Random Alignment variant performs the worst, as random alignment fails to guide the model effectively. This emphasizes the necessity of meaningful alignment strategies and demonstrates that our LLM-summarized reasoning chain provides a highly useful label.
Table 3: Generalization Analysis on different LLM and GNN backbones.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP
[…]
GT GNP 0.8704 0.7391 0.8880 0.7696
G-Retriever 0.8705 0.7366 0.9115 0.7739
Align-GRAG (Ours) 0.8992 0.7626 0.9241 0.7789
GAT GNP 0.9061 0.7291 0.8989 0.7676
G-Retriever 0.7960 0.7414 0.8953 0.7737
Align-GRAG (Ours) 0.9151 0.7309 0.9151 0.7573
GCN GNP 0.7545 0.7298 0.8682 0.7564
G-Retriever 0.8592 0.7352 0.9007 0.7521
Align-GRAG (Ours) 0.8574 0.7377 0.9152 0.7592
4.3 Generalization Analysis
In this section, we analyze whether the effectiveness of Align-GRAG can generalize across different GNN backbones and LLM sizes. Experiments are conducted on both ExplaGraphs and WebQSP datasets, with various GNN backbones, including GT (GraphTransformer) [37], GAT [38], and GCN [46], as well as LLM backbones of different sizes. As shown in Table 3, our method consistently achieves the best performance across most settings (9 out of 12 experimental settings), demonstrating its strong adaptability. Notably, our method shows the most significant improvement over GNP and G-Retriever when using GT as the backbone, highlighting that GT is particularly well-suited for learning dual alignment of node and representation. On the other hand, we observe that results with the -13B LLM generally outperform those with the -7B LLM, especially when GCN is used as the backbone. This suggests that larger LLMs play a crucial role in effectively learning and reasoning over graph structures.
4.4 Evaluation of Representation Alignment
@@
Figure 2: Representation Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
@@
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we calculate the Cosine Similarity scores between the graph embeddings and the language embeddings of the query, summarization, and textualized graph, respectively, on test set. The graph embeddings include both the unaligned embeddings and the aligned embeddings (optimized using our contrastive loss). From the results in Figure 2, we observe that as the alignment degree increases, Cosine Similarity scores improve. For instance, similarity with summarization embeddings improves significantly as alignment progresses from 40 to 80 steps. For query and textualized graph embeddings, scores rise sharply between 0-4 steps. This indicates that training effectively reduces the representation gap between graph and language embeddings. After 8 steps, aligned embeddings consistently achieve higher similarity scores then unaligned across all text representations (query, summarization, and textualized graph), demonstrating the effectiveness of our approach. However, as shown in experiment of Appendix E, the alignment degree does not necessarily improve indefinitely. When it reaches a certain level, excessive alignment may over-alter the original graph information, compromising its accuracy.
5 Related Work
Retrieval-Augmented Generation (RAG). RAG [13, 10, 13] has been extensively studied to address the challenges of hallucination and outdated information [8, 9], which often lead to untrustworthy outputs from LLMs [47]. By integrating information retrieval systems, RAG has demonstrated its effectiveness in real-world applications, enhancing reliability of generated responses [48, 49, 50, 51]. The RAG pipeline mainly includes: pre-retrieval (e.g., query rewriting and query expansion), retrieval, post-retrieval (e.g., reranking), and generation [13]. Among these, reranking has emerged as a key technique for refining retrieval results. Initial retrieval [52, 53] often relies on simple similarity scoring. Reranking [22, 21, 44], on the other hand, employs more sophisticated models to reassess and reorder the initially retrieved documents based on their relevance to the query, thereby improving the quality and relevance of retrieved documents. For instance, bge-reranker [21] leverages cross-encoders that perform full attention over the input pair, providing higher accuracy compared to embedding-based models. However, in GRAG scenarios, reranking may disrupt the inherent structural information of graph. Our proposed Align-GRAG effectively performs post-retrieval processing to extract relevant knowledge, while preserves the graph structural information.
Large Language Models on Graph. Graphs, composed of nodes and edges, are essential for modeling real-world relationships across various domains. In recent years, GNN [37, 38, 46, 54] have emerged as a powerful tool for encoding graph structures. With the rise of LLMs and their demonstrated impressive capabilities, there is growing interest in integrating LLMs with graph learning techniques to enhance graph-based tasks [55, 56, 57, 58]. Pioneering works [59, 31, 60] have incorporated GNNs by feeding graph tokens into LLMs. Beyond basic integration, recent studies [61, 62, 63] have focused on tighter architectural fusion, embedding graph neural layers within transformer architectures to enable seamless interaction between graph reasoning and natural language understanding . However, most existing researches [58] focus on tasks like node and graph classification. Our work addresses the challenges of retrieval-based graph QA, leveraging retrieval-augmented methods and graph learning techniques while overcoming the limitations of current frameworks.
Graph RAG. Traditional RAG often struggles to capture structured relational knowledge or global context. Graph Retrieval-Augmented Generation (GRAG) addresses these limitations by leveraging graph-structured knowledge for improved retrieval and reasoning [42, 41, 64, 65]. GRAG utilizes graph databases (e.g., Freebase [66], Wikidata [67]) to retrieve graph elements like triples, paths, or subgraphs [14, 68, 45, 69]. Some approaches [14, 15, 70, 71] focus on constructing large-scale graphs from text. For example, GraphRAG [14] uses LLMs to extract entities, partitions graphs using hierarchical clustering, and generates summaries to support retrieval. Other methods [72, 73, 18, 19] encode graph data with GNNs and integrate it with language models for enhanced reasoning, while some translate natural language queries into logical forms for knowledge graph retrieval [74, 75, 76, 27, 28]. More recently, LLMs have been used as iterative agents for reasoning over knowledge graphs [30, 26, 20, 29]. However, existing GRAG approaches primarily emphasize optimizing retrieval but often lack effective post-retrieval strategies, leading to shallow subgraph-query similarity. To address this, our Align-GRAG introduces a novel dual alignment mechanism that bridges retriever and generator, improving overall performance.
6 Conclusion
This work proposes a novel reasoning-guided dual alignment framework, designed to tackle two key challenges in GRAG: the retrieval of irrelevant nodes causing noisy inputs and the representation gap between graph structures and language models. Align-GRAG introduces an innovative graph aligner module for dual alignment of knowledge and representation, enabling the effective pruning of irrelevant nodes while unifying graph-language representations. Extensive experiments on the GraphQA benchmark demonstrate that Align-GRAG consistently outperforms strong baselines.
Appendix A Prize-Collecting Steiner Tree
To achieve the PCST algorithm, we assign ‘prize’ to nodes and edges based on their similarity to a given query. Relevance is determined through the ranked sets of cosine similarity, $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$, as follows:
$\text{prize}(n)=\begin{cases}k-i,&\text{if }n\in\mathcal{V}_{k}\text{ and }n\text{ is the }i\text{-th ranked node}\\ 0,&\text{otherwise}\end{cases}$ (12)
$\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}\left(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)-\text{cost}(S)\right)$ (13)
Appendix B Experimental Settings
Table 4: Statistics of GraphQA Benchmark [18].
Dataset ExplaGraphs SceneGraphs WebQSP
#Training 1,659 59,978 2,826
#Validation 553 19,997 245
#Test 554 20,025 1,628
B.1 Datasets and Metrics.
Following G-retriever [18], we conducted experiments on GraphQA benchmark [18], including ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 4, with a data split of train:validation:test = 6:2:2. The split for WebQSP is derived from the original dataset [77]. For ExplaGraphs and SceneGraphs, we use Accuracy as the evaluation metric, and for WebQSP, we employ F1, Hit@1, and Accuracy to evaluate model performance.
B.2 Baselines.
To evaluate the effectiveness of our proposed method, we compare it with four categories of baseline:
  • Inference-only. This category includes models that leverage frozen LLMs for question answering by using textual graph and query as input. Various prompt designs are employed, including: Zero-shot (answers questions directly based on retrieved information), Zero-CoT [35] (enhances zero-shot reasoning by appending the phrase "Let’s think step by step."), CoT-BAG [41] (adds "Let’s construct a graph with the nodes and edges first." after providing the textual description of graph), and KAPING [42] (retrieves relevant graph triples and prepends them to the input question).
  • Raw Fine-tuning. In this configuration, following [18], we fine-tune the LLM using parameter-efficient tuning methods, without incorporating advanced reranking techniques. It includes popular approaches like Prompt Tuning [43] and LoRA [39].
  • Reranker-based. These baselines use reranking models to refine the ranking of documents before inputting it into the LLM. Including: GTE (General Textual Embedding) [22], with two variants (gte-base, 109M parameters; gte-large, 335M parameters) developed by Alibaba DAMO Academy, trained on large-scale relevance text pairs to enhance retrieval accuracy. BGE (BAAI General Embedding) [21], with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), are cross-encoders optimized for retrieval-augmented LLMs, offering higher accuracy at the cost of efficiency. G-RAG and G-RAG-RL [44] use GNNs as rerankers, leveraging document connections and semantic information from abstract meaning representation graphs to improve context-aware ranking.
  • GNN-based. This category integrates GNN encoders with LLM embeddings. GraphToken [31] encodes graph structures as explicit prompts for LLMs, boosting graph reasoning performance. GNP [19] introduces a plug-and-play approach using a GNN encoder and cross-modality pooling to enhance LLMs with knowledge graphs. G-Retriever [18] establishes the GraphQA benchmark and employs a RAG method with soft prompting, improving graph-based QA, understanding, and mitigating hallucination. GRAG [45] introduces a divide-and-conquer strategy for efficient textual subgraph retrieval and incorporates both text and graph views into LLMs for graph context-aware generation.
B.3 Implementation Details.
For the retrieval process, we use the same retrieval results obtained through the PCST algorithm across all baselines to ensure a fair comparison. We use GraphTransformer [37], GAT [38], and GCN [46] as GNN encoders, and Llama-2-7b-hf and Llama-2-13b-hf [2] as generators. Both reranker-based and GNN-based methods apply LoRA for fine-tuning, except for GraphToken, whose results are from G-retriever [18]. All methods are compared using the same training hyperparameters where applicable, ensuring fair comparability (e.g., GNN layers and LoRA rank). In the Aligner module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 5. To summarize the reasoning chain, we employ Llama-3.1-70B-Instruct [78]. The models were trained on the training set, with optimal hyperparameters and early stopping determined using the validation set. Results are reported on the test set. All experiments were run on two NVIDIA A100 GPUs (80GB each). To implement baselines, for GNP [19], we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL [44], we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE [22] and BGE-reranker [21] using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model [45], we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach.
Appendix C Evaluating the Efficiency after Node Pruning

### v2

@@ 5.4. Main Results
Table 2. Performance comparison using the same retrieval settings for all methods. Llama-2-7b-hf is used as generator. The table reports the mean and standard deviation (std) results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report both F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from (He et al., 2024). The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement $(\triangle)$’ represents the gain of Align-GRAG over the second-best baseline. ‘*’ indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better.
@@
Inference-only
[…]
@@
[…]
@@
[…]
Improvement $(\triangle)$ +2.12%* +1.39% +1.68% +1.88%* +3.58%*
In this section, we conduct extensive experiments on the ExplaGraphs, SceneGraphs, and WebQSP datasets, comparing the performance of our method with 16 baseline methods. All baselines are implemented using the same retrieved text, and the same backbones, Llama2-7b (Touvron et al., 2023) and GraphTransformer (Shi et al., 2020). As shown in Table 2, the overall results demonstrate the effectiveness of Align-GRAG, which achieves performance improvements across all metrics on three datasets. Notably, on WebQSP dataset, Align-GRAG achieves a remarkable improvement in Accuracy compared to the second-best method. This improvement can be attributed to the innovative design of Align-GRAG, particularly its alignment-based pruning and optimization strategies for graph structures, showcasing its potential in complex graph reasoning and knowledge-based question answering tasks.
Comparing with Inference-only methods. Inference-only methods rely solely on the reasoning capabilities of LLMs without incorporating task-specific optimization, which often leads to subpar performance. Specifically, the Zero-shot approach, which does not depend on any carefully constructed prompts, suffers from limited performance due to its inability to leverage task-specific knowledge. In contrast, ToG achieves superior performance by iteratively executing beam search on KG, enabling the discovery of promising reasoning paths. These results underscore ToG’s effectiveness within inference-only approaches.
@@
Comparing with reranker-based methods. gte-large achieves the second-best performance on both the ExplaGraphs and WebQSP datasets (F1 metric), outperforming other reranker approaches. However, it still falls significantly short of Align-GRAG. This demonstrates that simply performing node and triple reranking is insufficient to fully capture the graph structural information. Although G-RAG-RL also utilizes a GNN encoder to model graph structures, it relies on similarity with the query as a label, resulting in poor performance. This validates the effectiveness of our approach, which uses reasoning chains summarized by LLM as an optimization target. By aligning graph reasoning with LLM-derived reasoning chains, Align-GRAG achieves superior performance, showing that this alignment is crucial for graph-based reasoning tasks.
@@
5.5. Ablation Study
Figure 3. Ablation study with different alignment strategy.
We perform an ablation study on WebQSP using three metrics to evaluate how the proposed modules contribute to performance improvement. As illustrated in Figure 3, we examine the following variants: (1) w/o Node Alignment: Removes the KL loss and knowledge pruning components. (2) w/o Representation Alignment: Excludes the contrastive loss and the graph embedding input to the LLM. (3) w/o Both: Removes both Node Alignment and Representation Alignment. (4) Random Alignment: Optimizes the aligner module using randomly generated labels. We observe that removing the Node Alignment module significantly lowers all evaluation metrics compared to the full model (Align-GRAG), indicating the effectiveness of node-level alignment optimization and node pruning. In contrast, while excluding the Representation Alignment module also leads to a performance decline, the impact is comparatively less significant. Performance further deteriorates when both modules are removed (w/o Both), highlighting the critical role of dual alignment between node and representation in filtering irrelevant knowledge and bridging the representation gap between the graph and language. The Random Alignment variant performs the worst, as random alignment fails to guide the model effectively. This emphasizes the necessity of meaningful alignment strategies and demonstrates that our LLM-summarized reasoning chain provides a highly useful label.
5.6. Generalization Analysis
Table 3. Generalization Analysis on different LLM and GNN backbones.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf Llama-3.1-8B
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP ExplaGraphs WebQSP
@@ 5.6. Generalization Analysis
GraphTransformer GNP 0.8704 0.7391 0.8880 0.7696 0.8892 0.7557
G-Retriever 0.8705 0.7366 0.9115 0.7739 0.9014 0.7482
Align-GRAG 0.8992 0.7626 0.9241 0.7789 0.9321 0.7712
GAT GNP 0.9061 0.7291 0.8989 0.7676 0.9062 0.7521
G-Retriever 0.7960 0.7414 0.8953 0.7737 0.9111 0.7719
Align-GRAG 0.9151 0.7309 0.9151 0.7573 0.9214 0.7742
GCN GNP 0.7545 0.7298 0.8682 0.7564 0.8572 0.7495
G-Retriever 0.8592 0.7352 0.9007 0.7521 0.8512 0.7527
Align-GRAG 0.8574 0.7377 0.9152 0.7592 0.8596 0.7573
In this section, we evaluate whether Align-GRAG generalizes across both GNN architectures and LLM scales on ExplaGraphs (Accuracy) and WebQSP (Hit@1) datasets. Table 3 summarizes results for three GNN backbones—GT (GraphTransformer) (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016)—and three LLM families/sizes (Llama-2-7b-hf, Llama-2-13b-hf, Llama-3.1-8B), comparing against GNP and G-Retriever. We can observe that Align-GRAG attains the best score in 15/18 configurations, indicating robust generalization to diverse encoder–reasoner pairings.
What is more, we can find the strongest gains with GraphTransformer (GT). Using GT as the graph encoder yields the largest margins over GNP and G-Retriever on both datasets and across LLM sizes. This suggests GT’s capacity to model long-range dependencies synergizes with Align-GRAG’s dual alignment of nodes and representations, amplifying retrieval and reasoning quality.
Besides, benefits persist across attention-based GNNs. With GAT, Align-GRAG remains consistently competitive and often best, reinforcing that alignment mechanisms complement attention-driven neighborhood aggregation. Larger LLMs help, especially with GCN. Performance with 13B models generally surpasses 7B counterparts; the gap is most pronounced under the GCN backbone, where the LLM must compensate for weaker expressivity in the graph encoder. This indicates that larger LLMs better exploit aligned graph signals for multi-hop reasoning.
On the other hand, improvements hold on both ExplaGraphs and WebQSP datasets, demonstrating that Align-GRAG transfers well from explanation-centric graph tasks to question answering over knowledge graphs. Overall, Align-GRAG is architecture-agnostic yet particularly effective when paired with more expressive GNNs and larger LLMs, underscoring the importance of both high-capacity graph encoders and language models for learning fine-grained node-representation alignment.
5.7. Evaluation of Representation Alignment
@@
Figure 4. Representation Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
@@
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we calculate the Cosine Similarity scores between the graph embeddings and the language embeddings of the query, summarization, and textualized graph, respectively, on test set. The graph embeddings include both the unaligned embeddings and the aligned embeddings (optimized using our contrastive loss). From the results in Figure 4, we observe that as the alignment degree increases, Cosine Similarity scores improve. For instance, similarity with summarization embeddings improves significantly as alignment progresses from 40 to 80 steps. For query and textualized graph embeddings, scores rise sharply between 0-4 steps. This indicates that training effectively reduces the representation gap between graph and language embeddings. After 8 steps, aligned embeddings consistently achieve higher similarity scores than unaligned across all text representations (query, summarization, and textualized graph), demonstrating the effectiveness of our approach. However, as shown in section 5.9.2, the alignment degree does not necessarily improve indefinitely. When it reaches a certain level, excessive alignment may over-alter the original graph information, compromising its accuracy.
5.8. Evaluating the Efficiency of Node Pruning